OpenStudy (anonymous):

A recipe calls for 10 cups of oatmeal. Benjamin wants to make fewer servings than the recipe makes. Which expression represents how many cups of oatmeal Benjamin should use A. 0.8 10 B. 1.8 10 C. 0.2 10 D. 1.2 10

OpenStudy (love_ranaa):

1/5 fewer servings of 10 is: 10 - (1/5)100 = 10(1 - 1/5) = 10(1 - 0.2) = 10*0.8 = 0.8 x 10 the answer is A. 0.8 x 10.